OpenStudy (anonymous):

find the equation of the line containing given the pair of points (-4,0) and (0,1)

OpenStudy (anonymous):

y - y1 = m(x -x1) 0 - 1 = m(-4 - 0) -1 = m(-4) m = 4 y = mx + b 1 = 4(0) + b 1 = 0 + b b = 1 y = 4x + 1

OpenStudy (anonymous):

You can use the slope formula, then point slope